Write an equations in slope-intercept form of the line satisfying the given conditions.
Passes through (5,4) and (6,-1)
and
slope= 1/3 and passes through (-2,8)
and last question of the day, parralel to the graph of 9x+ 3y=6.
Please email me if you get the answer to any of these. Thanks again.

Found 2 solutions by checkley71, zman0652:
Answer by checkley71(8403)

SLOPE IS (-1-4)/(6-5) OR 15/1 OR -5 THEN Y=-5X+B THUS THE Y INTERCEPT IS
(5*5)+4 OR 29 OR (6*5)-1 OR 29 THUS THE EQUATION IS Y=5X+29
Y=-1/3X+B OR THE Y INTERCEPT IS (2*1/3)+8 OR 2/3+8 OR 8 2/3 OR Y=1/3X+8 2/3

Answer by zman0652(1)